# EthioLLM: Multilingual Large Language Models for Ethiopian Languages with Task Evaluation

### v2

@@ Abstract
[…]
@@ 3.1.   Training Data and Languages
Amharic (amh) Ge’ez Afro-Asiatic / Ethio-Semitic 57M yes $\star,\dagger,\ast$⋆ , † , ∗
153,509,645
9,365,829
English (eng) Latin Indo-European / Germanic 1268M yes $\star,\ast$⋆ , ∗
76,587,128
2,275,996
Afaan Oromo (orm) Latin Afro-Asiatic / Cushitic 37M yes $\dagger,\star,\ast$† , ⋆ , ∗
22,448,422
1,040,175
Ge’ez (gez) Ge’ez Afro-Asiatic / Ethio-Semitic UNK no $\dagger$
1,086,578
95,899
Somali (som) Latin Afro-Asiatic / Cushitic 22.3M no $\star,\ast$⋆ , ∗
17,589,974
558,161
Tigrinya (tir) Ge’ez Afro-Asiatic / Ethio-Semitic 9M yes $\dagger,\star,\ast$† , ⋆ , ∗
28,290,680
1,344,586
@@ 4.2.   Machine Translation (MT)
MT is a widely used NLP application that automatically translates one language to another to facilitate communication between people who speak different languages Forcada (2017). Many machine translation works Biadgligne and Smaïli (2021); Gezmu et al. (2021); Teshome et al. (2015); Abate et al. (2018); Teshome and Besacier (2012); Ashengo et al. (2021); Ambaye and Yared (2000) use different statistical machine translation approaches. The work by Belay et al. (2022) is done by fine-tuning an available multilingual pre-trained model (M2M100 418M) from NLLB Team et al. (2022). Most of the works use traditional approaches and cover two parallel languages, except for the work of Abate et al. (2018), which covers English and five Ethiopian languages (Amharic, Tigrinya, Afan-Oromo, Wolaytta, and Ge’ez). We combined available MT datasets from the works of Biadgligne and Smaïli (2021); Abate et al. (2018); Gezmu et al. (2021); Belay et al. (2022) and HornMT online repository [^7] https://github.com/asmelashteka/HornMT/tree/main into one new benchmark dataset. We present the statistics of the new MT benchmark dataset in Table 2.
@@ 4.3.   Hate Speech
Detecting hate speech plays a crucial role in content moderation by identifying and screening out harmful or offensive language from online platforms, thereby fostering a safer online environment Davidson et al. (2017); Mathew et al. (2021). Detecting hate speech in low-resource languages is challenging due to sparse data, linguistic diversity, and complex cultural nuances, making it difficult to develop accurate and contextually aware models Ousidhoum et al. (2019); Ayele et al. (2023).
@@
Sentiment analysis constitutes a prominent domain in the field of Natural Language Processing, focusing on the automated detection of emotions or opinions expressed in digital content, including social media posts, blog articles, and reviews. This discipline leverages computational techniques to discern and classify the sentiments or viewpoints encoded in textual data sourced from the internet Agarwal et al. (2011); Taboada et al. (2011).
@@ 4.6.   Part-of-Speech (POS) Tagging
POS tagging stands as one of the sequence labeling tasks within the realm of NLP, where each word (token) in a given sentence is assigned a part of speech tag or another philological class Keiper et al. (2016). To assess the POS tagging capabilities of our models, we employed a publicly available Amharic POS tagging dataset comprising 33,940 sentences (440,941 words) from the research of Gashaw and Shashirekha (2020) and data from the Habit project [^8] https://habit-project.eu/wiki/SetOfEthiopianWebCorpora for Amharic, Tigrinya, Oromo, and Somali, with the Habit project data yet to be evaluated by researchers. We merged two Amharic datasets to create a novel benchmark dataset. Additionally, we curated a new Ge’ez POS tagging test dataset to evaluate the zero-shot performance of our models. The statistics of this new benchmark dataset are presented in Table 2.

### v3

@@ Abstract
Atnafu Lambebo Tonja1,2,∗,†, Israel Abebe Azime3,∗,†, Tadesse Destaw Belay1,†,
Mesay Gemeda Yigezu1,†, Moges Ahmed Mehamed4,†, Abinew Ali Ayele5,6,†,
Ebrahim Chekol Jibril7,†, Michael Melese Woldeyohannis8,†, Olga Kolesnikova1,
Philipp Slusallek2, Dietrich Klakow2, Shengwu Xiong3,
Seid Muhie Yimam6,†
Ethio NLP, 1 Instituto Politécnico Nacional, Mexico, 2 Lelapa AI, 3 Saarland University, Germany,
4 Wuhan University of Technology, China, 5 Bahir Dar University, Ethiopia, 6 Universität Hamburg, Germany,
7 Istanbul Technical University, Turkey, 8 Addis Ababa University, Ethiopia
@@ 3.1.   Training Data and Languages
Amharic (amh) Ge’ez Afro-Asiatic / Ethio-Semitic 57M yes $\star,\dagger,\ast$⋆ , † , ∗ 153,509,645 9,365,829
English (eng) Latin Indo-European / Germanic 1268M yes $\star,\ast$⋆ , ∗ 76,587,128 2,275,996
Afaan Oromo (orm) Latin Afro-Asiatic / Cushitic 37M yes $\dagger,\star,\ast$† , ⋆ , ∗ 22,448,422 1,040,175
Ge’ez (gez) Ge’ez Afro-Asiatic / Ethio-Semitic UNK no $\dagger$ 1,086,578 95,899
Somali (som) Latin Afro-Asiatic / Cushitic 22.3M no $\star,\ast$⋆ , ∗ 17,589,974 558,161
Tigrinya (tir) Ge’ez Afro-Asiatic / Ethio-Semitic 9M yes $\dagger,\star,\ast$† , ⋆ , ∗ 28,290,680 1,344,586
@@ 4.2.   Machine Translation (MT)
MT is a widely used NLP application that automatically translates one language to another to facilitate communication between people who speak different languages Forcada (2017). Many machine translation works Biadgligne and Smaïli (2021); Gezmu et al. (2021); Teshome et al. (2015); Abate et al. (2018); Teshome and Besacier (2012); Ashengo et al. (2021); Ambaye and Yared (2000) use different statistical machine translation approaches. The work by Belay et al. (2022) is done by fine-tuning an available multilingual pre-trained model (M2M100 418M) from NLLB Team et al. (2022). Most of the works use traditional approaches and cover two parallel languages, except for the work of Abate et al. (2018), which covers English and five Ethiopian languages (Amharic, Tigrinya, Afan-Oromo, Wolaytta, and Ge’ez). We combined available MT datasets from the works of Biadgligne and Smaïli (2021); Abate et al. (2018); Gezmu et al. (2021); Belay et al. (2022) and HornMT online repository [^7] https://github.com/asmelashteka/HornMT into one new benchmark dataset. We present the statistics of the new MT benchmark dataset in Table 2.
@@ 4.3.   Hate Speech
Detecting hate speech plays a crucial role in content moderation by identifying and screening out harmful or offensive language from online platforms, thereby fostering a safer online environment Davidson et al. (2017); Mathew et al. (2021). Detecting hate speech in low-resource languages is challenging due to sparse data, linguistic diversity, and complex cultural nuances, making it difficult to develop accurate and contextually aware models ayelechallenges; Ousidhoum et al. (2019); muhie2019analysis.
@@
Sentiment analysis constitutes a prominent domain in the field of Natural Language Processing, focusing on the automated detection of emotions or opinions expressed in digital content, including social media posts, blog articles, and reviews. This discipline leverages computational techniques to discern and classify the sentiments or viewpoints encoded in textual data sourced from the internet Agarwal et al. (2011); Taboada et al. (2011); yimam-etal-2020-exploring.
@@ 4.6.   Part-of-Speech (POS) Tagging
POS tagging stands as one of the sequence labeling tasks within the realm of NLP, where each word (token) in a given sentence is assigned a part of speech tag or another philological class Keiper et al. (2016). To assess the POS tagging capabilities of our models, we employed a publicly available Amharic POS tagging dataset comprising 33,940 sentences (440,941 words) from the research of Gashaw and Shashirekha (2020) and data from the Habit project [^8] https://habit-project.eu/wiki for Amharic, Tigrinya, Oromo, and Somali, with the Habit project data yet to be evaluated by researchers. We merged two Amharic datasets to create a novel benchmark dataset. Additionally, we curated a new Ge’ez POS tagging test dataset to evaluate the zero-shot performance of our models. The statistics of this new benchmark dataset are presented in Table 2.